### v2

# Nonlinear transport theory at the order of quantum metric

Abstract
Quantum metric, a probe to spacetime of the Hilbert space, has been found measurable in the nonlinear electronic transport and attracted tremendous interest. We show that the quantum metric is only a tip of the iceberg, by deriving unknown 11 out of 14 formulas of the quadratic nonlinear conductivity under the same symmetry that the quantum metric emerges. The formulas allow us to determine nonzero nonlinear conductivity elements for the magnetic point groups, calculate the quadratic nonlinear conductivity for arbitrary Hamiltonian, e.g., the even-layered MnBi2Ti4 thin films with and without the $C_{3}$ symmetry, and formulate scaling laws at the order of quantum metric to distinguish different mechanisms in the recent experiments. The theory is a comprehensive description of the quadratic nonlinear transport and will facilitate future experiments and applications.
Introduction.– The nonlinear Hall effect is characterized by a transverse voltage nonlinearly dependent on the longitudinal driving current. It has received tremendous attention because of its potential as a measurement method and for device applications Sodemann and Fu (2015); Low et al. (2015); Ma et al. (2019); Du et al. (2018); Kang et al. (2019); Du et al. (2019); Nandy and Sodemann (2019); Xiao et al. (2019); Isobe et al. (2020); Du et al. (2021a, b); Kumar et al. (2021); Ho et al. (2021); He et al. (2021); Watanabe and Yanase (2020); Tiwari et al. (2021); König and Levchenko (2021); Duan et al. (2022); He et al. (2022); Huang et al. (2022); Kaplan et al. (2023); Min et al. (2023); Kaplan et al. (2024); Chen et al. (2024a); Wang et al. (2024); Xiang and Wang (2024); Chen et al. (2024b); Yao et al. (2024); He et al. (2024). More excitingly, without time-reversal and inversion symmetries but under a combined time-reversal and inversion symmetry, the nonlinear Hall effect can be contributed by the quantum metric Gao et al. (2014); Wang et al. (2021); Liu et al. (2021); Gao et al. (2023); Wang et al. (2023); Fang et al. (2024), the real part of quantum geometry Provost and Vallee (1980); Resta (2011); Törmä et al. (2022); Törmä (2023); Hetényi and Lévay (2023); Bouhon et al. (2023), which measures the distance between quantum states, thus can provide spacetime insight to the challenges in condensed matter Liu et al. (2024). The nonlinear Hall effect from the quantum metric dipole ($\chi^{\mathrm{QMD}}$ and $\mathcal{G}_{l}^{ab}$ in Tab. 1) is essentially an anomalous Hall effect but with the driving current acting twice. In this sense, the disorder contributions, such as the side-jump and skew-scattering Nagaosa et al. (2010), cannot be ignored Du et al. (2019), inspiring an increasing interest in exhausting all contributions to the nonlinear transport Ma et al. (2023); Atencia et al. (2023).
Figure 1: Hierarchy of the quadratic nonlinear transport, including (i) the Drude mechanism in absence of time reversal and inversion symmetries, (ii) the Berry curvature dipole Sodemann and Fu (2015); Low et al. (2015), side-jump, and skew-scattering mechanisms Du et al. (2019, 2021b) under time-reversal but without inversion symmetry, and (iii) the quantum metric dipole (Berry connection polarizability) Gao et al. (2014); Wang et al. (2021); Liu et al. (2021) and mixed Berry and skew-scattering contributions Ma et al. (2023), as well as 2nd-order side-jump, 2nd-order skew-scattering, mixed Berry-side-jump, and mixed side-jump-skew-scattering revealed in this work, in absence of time-reversal and inversion symmetries but under a combined time-reversal and inversion symmetry. They can also be roughly categorised by the power law of the scattering time $\tau$. The color bar indicates the Berry curvature $\mathbf{\Omega}$.
Table 1: Formulas of the quadratic nonlinear conductivity at the order of quantum metric, including the quantum metric dipole (QMD), Drude, mixed Berry curvature and skew-scattering (BSK), mixed Berry curvature and side-jump (BSJ), second-order side-jump (2SJ) and skew-scattering (2SK), mixed side-jump and skew-scattering (SSK). $\tau$ is the scattering time, $-e$ is the electron charge, $\varepsilon^{abc}$ is the anti-symmetric tensor, Fermi distribution $f_{l}^{(0)}=1/[\exp[(\epsilon_{l}-\epsilon_{F})/k_{B}T]+1]$ with the Fermi energy $\epsilon_{F}$, $l$ stands for the band index $\gamma$ and wave vector $\mathbf{k}=|k|\left(\cos\phi,\sin\phi\right)$ in 2D, $\Omega^{a}_{l}$ is the Berry curvature Xiao et al. (2010), $\mathcal{G}^{ab}_{l}$ is the Berry connection polarizability tensor, which is related to the quantum metric tensor $\mathrm{Re}(A^{\gamma\gamma^{\prime}}_{a}A^{\gamma^{\prime}\gamma}_{b})$ for two-band cases Gao et al. (2014); Wang et al. (2021); Liu et al. (2021), Berry connection $A^{\gamma\gamma^{\prime}}_{a}\equiv i\langle\gamma|\partial_{\mathbf{k}}^{a}|\gamma^{\prime}\rangle$ Xiao et al. (2010), $\partial^{b}_{\mathbf{k}}$$\equiv$$\partial/\partial k_{b}$, $\hat{\mathcal{H}}$ is the Hamiltonian, $\epsilon_{l}$ is the eigen energy, group velocity $v_{l}^{a}$ = $(1/\hbar)\partial\epsilon_{l}/\partial k_{a}$, $\partial_{\epsilon_{l}}\equiv\partial/\partial\epsilon_{l}$, side-jump velocity $v_{l}^{sj,a}=\sum_{l^{\prime}}\varpi^{S}_{ll^{\prime}}\delta r^{a}_{l^{\prime}l}$, $\varpi^{S}_{ll^{\prime}}$ ($\varpi_{l^{\prime}l}^{A}$) is the symmetric (anti-symmetric) scattering rate between states $l$ and $l^{\prime}$ (Sec. SIB of Sup ), $\delta r^{a}_{ll^{\prime}}$ is the coordinates shift with $V_{ll^{\prime}}$=$\langle l|\hat{V}_{imp}|l^{\prime}\rangle$ Du et al. (2019); Sinitsyn (2008), where the $\delta$-correlated random potential $\hat{V}_{imp}(\mathbf{r})$=$\sum_{i}V_{i}\delta(\mathbf{r}-\mathbf{R}_{i})$ can give the Gaussian $\langle V^{2}_{i}\rangle_{dis}$=$V^{2}_{0}$ and non-Gaussian correlations $\langle V^{3}_{i}\rangle_{dis}$=$V^{3}_{1}$ Du et al. (2019, 2021b); Sinitsyn (2008), leading to the Gaussian (intrinsic skew-scattering) $\varpi_{l^{\prime}l}^{(4A)}$ and non-Gaussian (extrinsic skew-scattering) $\varpi_{l^{\prime}l}^{(3A)}$ parts to $\varpi_{l^{\prime}l}^{A}$. $\bar{O}_{ll^{\prime}}^{a}$ ($\tilde{O}_{ll^{\prime}}^{a}$) is the symmetric (anti-symmetric) part of $O^{a}_{ll^{\prime}}$ Du et al. (2019), and $T_{ll^{\prime}}$ is the T-matrix Mahan (1990) (details in Secs. SI and SII of Sup ).
Known (3 out of 14) This work (11 out of 14)
$\chi_{2}^{\mathrm{BSK}}=\frac{\tau^{2}e^{3}}{2\hbar^{2}}\sum_{ll^{\prime}}{\varepsilon^{acd}\Omega_{l^{\prime}}^{d}}\varpi_{ll^{\prime}}^{(4A)}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
$\chi^{\mathrm{QMD}}=e^{2}\sum_{l}{\left(v_{l}^{a}\mathcal{G}_{l}^{bc}-v_{l}^{b}\mathcal{G}_{l}^{ac}\right)}\partial_{\epsilon_{l}}f^{(0)}_{l}$ Gao et al. (2014); Wang et al. (2021); Liu et al. (2021), $\chi^{\mathrm{BSJ}}=\frac{\tau e^{3}}{2\hbar}\sum_{l}{\varepsilon^{acd}\Omega_{l}^{d}v_{l}^{sj,b}\partial_{\epsilon_{l}}f_{l}^{(0)}}$,
$\chi^{\mathrm{Drude}}=-\frac{\tau^{2}e^{3}}{2\hbar^{2}}\sum_{l}v_{l}^{a}\partial_{\mathbf{k}}^{b}\partial_{\mathbf{k}}^{c}f_{l}^{(0)}$ Sodemann and Fu (2015), $\chi_{1}^{2\mathrm{SK}}=\frac{e^{3}\tau^{4}}{2\hbar^{2}}\sum_{ll^{\prime}l^{\prime\prime}}\left(\partial_{\mathbf{k}^{\prime\prime}}^{c}v_{l^{\prime\prime}}^{a}\right)\varpi_{l^{\prime}l}^{A}\varpi_{l^{\prime\prime}l^{\prime}}^{A}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
$\chi^{\mathrm{BSK}}_{1}=\frac{\tau^{2}e^{3}}{2\hbar^{2}}\sum_{ll^{\prime}}{\varepsilon^{acd}\Omega_{l^{\prime}}^{d}}\varpi_{ll^{\prime}}^{(3A)}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$ Ma et al. (2023), $\chi_{2}^{2\mathrm{SK}}=\frac{e^{3}\tau^{4}}{2\hbar^{2}}\sum_{ll^{\prime}l^{\prime\prime}}v_{l^{\prime\prime}}^{a}\left(\partial_{\mathbf{k}}^{c}-\partial_{\mathbf{k}^{\prime}}^{c}\right)\varpi_{l^{\prime}l}^{A}\varpi_{l^{\prime\prime}l^{\prime}}^{A}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
$\chi_{1}^{2\mathrm{SJ}}=\frac{\tau^{2}e^{3}}{2}\sum_{ll^{\prime}}\left(v_{l^{\prime}}^{a}+v_{l}^{a}\right)\bar{O}_{l^{\prime}l}^{c}v_{l}^{sj,b}\partial_{\epsilon_{l}}f_{l}^{(0)}$,
$\chi_{2}^{2\mathrm{SJ}}=\frac{\tau^{2}e^{3}}{2\hbar}\sum_{ll^{\prime}}{(v_{l}^{sj,a}+v_{l^{\prime}}^{sj,a})\bar{O}_{ll^{\prime}}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}}$,
where $\chi_{3}^{2\mathrm{SJ}}=-\frac{\tau^{2}e^{3}}{2\hbar}\sum_{l}{\partial_{\mathbf{k}}^{b}v_{l}^{sj,a}}v_{l}^{sj,c}\partial_{\epsilon_{l}}f_{l}^{(0)}$,
$\mathcal{G}^{ab}_{l}\equiv 2e\mathrm{Re}\sum_{\gamma^{\prime}\neq\gamma}A^{\gamma\gamma^{\prime}}_{a}A^{\gamma^{\prime}\gamma}_{b}/(\epsilon^{\gamma}_{\mathbf{k}}-\epsilon^{\gamma^{\prime}}_{\mathbf{k}})$, $\chi_{1}^{\mathrm{SSK}}=\frac{\tau^{2}e^{3}}{2\hbar}\sum_{ll^{\prime}}\left(v_{l}^{a}+v_{l^{\prime}}^{a}\right)\tilde{O}_{l^{\prime}l}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
$\Omega^{a}_{l}\equiv$ $-2\varepsilon^{abc}\sum_{\gamma^{\prime}\neq\gamma}$ $\mathrm{Im}\langle\gamma|\partial^{b}_{\mathbf{k}}\hat{\mathcal{H}}|\gamma^{\prime}\rangle$ $\langle\gamma^{\prime}|\partial^{c}_{\mathbf{k}}\hat{\mathcal{H}}|\gamma\rangle$/$(\epsilon^{\gamma}_{\mathbf{k}}-\epsilon^{\gamma^{\prime}}_{\mathbf{k}})^{2}$, $\chi_{2}^{\mathrm{SSK}}=\frac{\tau^{3}e^{3}}{2\hbar}\sum_{ll^{\prime}l^{\prime\prime}}\left(v_{l^{\prime\prime}}^{a}+v_{l^{\prime}}^{a}\right)\varpi_{ll^{\prime}}^{A}\bar{O}_{l^{\prime\prime}l^{\prime}}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
$\delta r^{a}_{ll^{\prime}}\equiv i\langle l|\partial^{a}_{\mathbf{k}}|l\rangle$$-i\langle l^{\prime}|\partial^{a}_{\mathbf{k^{\prime}}}|l^{\prime}\rangle$$-(\partial^{a}_{\mathbf{k}}$+$\partial^{a}_{\mathbf{k^{\prime}}})\arg(V_{ll^{\prime}})$, $\chi_{3}^{\mathrm{SSK}}=\frac{\tau^{3}e^{3}}{2\hbar}\sum_{ll^{\prime}l^{\prime\prime}}\left(\varpi_{l^{\prime}l}^{A}+\varpi_{l^{\prime}l^{\prime\prime}}^{A}\right)v_{l^{\prime}}^{a}\bar{O}_{ll^{\prime}}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$
$O^{a}_{ll^{\prime}}$$\equiv$$(2\pi/\hbar)|T_{ll^{\prime}}|^{2}\delta r^{a}_{ll^{\prime}}\partial_{\epsilon_{l}}\delta(\epsilon_{l}-\epsilon_{l^{\prime}})$, $\quad\quad\quad+\frac{\tau^{3}e^{3}}{2\hbar}\sum_{ll^{\prime}}v_{l^{\prime}}^{a}v_{l}^{sj,b}\left(\partial_{\mathbf{k}}^{c}-\partial_{\mathbf{k}^{\prime}}^{c}\right)\varpi_{l^{\prime}l}^{A}\partial_{\epsilon_{l}}f_{l}^{(0)}$,
$1/\tau\approx 1/\tau_{\bf k}\equiv\sum_{\bf k^{\prime}}\varpi_{\mathbf{k}^{\prime}\mathbf{k}^{\prime}}^{(2)}\left[1-\cos\left(\phi-\phi^{\prime}\right)\right]\delta\left(\epsilon_{\mathrm{F}}-\epsilon_{l}\right)$. $\chi_{4}^{\mathrm{SSK}}=\frac{\tau^{3}e^{3}}{2\hbar^{2}}\sum_{ll^{\prime}}v_{l^{\prime}}^{sj,a}\left(\partial_{\mathbf{k}}^{c}-\partial_{\mathbf{k}^{\prime}}^{c}\right)\varpi_{l^{\prime}l}^{A}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$.
In this Letter, we try to formulate all the contributions to the quadratic nonlinear conductivities under the same symmetry that the quantum metric emerges, by treating the geometry and disorder on the same footing. We find 11 out of 14 conductivity formulas unknown in the literature. The formulas allow a symmetry analysis to determine nonzero nonlinear conductivity elements for the magnetic point groups at the order of quantum metric. By applying the formulas to even-layered MnBi2Te4 thin films as those in the experiments Gao et al. (2023); Wang et al. (2023), we find that the disorder-induced contributions could be comparable with the quantum metric part, even stronger by orders of magnitude in the clean limit. The formulas also help to extend the scaling law to the order of quantum metric, that is, the relation between linear longitudinal and nonlinear transverse resistivities, which is crucial for distinguishing mechanisms contributing to the nonlinear transport. This theory is not only at the order of quantum metric but a comprehensive description of the quadratic nonlinear transport under various symmetries.
Quadratic conductivity formulas at the order of quantum metric.– Experimentally, the nonlinear transport is measured as a double-frequency longitudinal or transverse voltage, in response to low frequency (10-1000 Hz) driving currents. Theoretically, it is equivalent to calculate double-frequency current density
$J_{a}(2\omega)=\chi_{abc}E_{b}(\omega)E_{c}(\omega),$ (1)
in response to low-frequency ($\omega\tau\ll 1$) electric fields $E_{b},E_{c}$, where $a,b,c\in\{x,y,z\}$. To treat the geometry and disorder contributions on the same footing, we adopt the semiclassical theory Du et al. (2019); Sinitsyn (2008) to calculate the nonlinear conductivity $\chi_{abc}$ (Secs. SI and SII of Sup ), in which each contribution has a clear physical origin, as shown by the hierarchy in Fig. 1. The calculation starts with the uniform Boltzmann equation $\partial f_{l}/\partial t+\dot{\mathbf{k}}\cdot\partial f_{l}/\partial\mathbf{k}=\mathcal{I}_{el}\{f_{l}\}$ of the distribution function $f_{l}$, where $\mathcal{I}_{el}\{f_{l}\}$ accounts for the elastic disorder scattering, $l$ stands for the band index $\gamma$ and wave vector $\mathbf{k}$, $\dot{\mathbf{k}}=-(e/\hbar)\mathbf{E}$, and $\mathbf{E}$ is the electric field. By decoupling the distribution function and scattering according to different mechanisms ($in$ for disorder-irrelavent, $sj$ and $sk$ for side-jump and skew-scattering, respectively)
$f_{l}$ $=$ $f_{l}^{in}+\delta f_{l}^{sj}+\delta f_{l}^{sk}+\delta f_{l}^{2sj}+\delta f_{l}^{2sk}+\delta f_{l}^{sk,sj},$
$\mathcal{I}_{el}\{f_{l}\}$ $=$ $\mathcal{I}^{in}+\mathcal{I}^{sj}+\mathcal{I}^{sk}+\mathcal{I}^{sk,sj},$ (2)
the distribution function can be found up to the second order of the electric field, where $\delta f_{l}^{2sj}+\delta f_{l}^{2sk}+\delta f_{l}^{sk,sj}$ and $\mathcal{I}^{sk,sj}$ are crucial for new findings. Finally, the nonlinear conductivity $\chi_{abc}$ can be extracted from the electric current $\mathbf{J}(\mathbf{E})=-e\sum_{l}\dot{\mathbf{r}}_{l}f_{l}$, where the velocity $\dot{\mathbf{r}}_{l}=\mathbf{v}_{l}-\dot{\mathbf{k}}\times\bm{\Omega}_{l}+\dot{\mathbf{k}}\times\nabla_{\mathbf{k}}\times\bm{\mathcal{G}}\mathbf{E}+\mathbf{v}_{l}^{sj}$ includes the components from the group velocity $\mathbf{v}_{l}$, Berry curvature $\bm{\Omega}$ Xiao et al. (2010), Berry connection polarizability $\bm{\mathcal{G}}$ (related to the quantum metric)Gao et al. (2014); Wang et al. (2021); Liu et al. (2021), and side-jump velocity $\mathbf{v}^{sj}$.
At the order of quantum metric, we find that the quadratic conductivity $\chi_{abc}$ can be decomposed into
$\chi_{abc}$ $=$ $\sum_{i=1}^{2}\chi_{i}^{\mathrm{2SK}}+\sum_{i=1}^{3}\chi_{i}^{\mathrm{2SJ}}+\sum_{i=1}^{4}\chi_{i}^{\mathrm{SSK}}+\chi^{\mathrm{BSJ}}+\chi_{2}^{\mathrm{BSK}}$ (3)
$+\chi_{1}^{\mathrm{BSK}}+\chi^{\mathrm{Drude}}+\chi^{\mathrm{QMD}},$
where the first row is our finding, including 11 formulas at the order of quantum metric, as compared in Tab. 1 with 3 known contributions on the second row. The first row can survive the same symmetry that the quantum metric dipole $\chi^{\mathrm{QMD}}$ emerges in the nonlinear transport (see Sec. SIIIA of Sup ), i.e., no time-reversal and inversion symmetries but under a combined time-reversal and inversion symmetry. Alternatively, the calculation can be based on the density matrix Atencia et al. (2022, 2023); Liu et al. (2023a); Mehraeen (2024), but does not give formulas as those in Tab. 1. Below, we show that the explicit form of the formulas is crucial.
Table 2: Nonzero quadratic conductivity elements $\chi_{abb}$ on the $x$-$y$ plane (defined as $J_{a}$ = $\chi_{abb}E_{b}^{2}$, $\{a,b\}\in\{x,y\}$, $abb$ is short for $\chi_{abb}$) at the order of quantum metric (all formulas in Tab. 1), can be found for only 9 out of the 122 magnetic point groups. The quantum metric dipole (QMD) Gao et al. (2014); Wang et al. (2021); Liu et al. (2021) as well as BSJ and BSK here have the anti-symmetric constraint $\chi_{abc}$ = $-\chi_{cba}$ Du et al. (2021b), so they have less nonzero elements Zhang et al. (2023); Zhu et al. (2024) than 2SJ, 2SK, and SSK found in this work. The $C_{2}$ axis is assumed in order as the $y$- and $x$- axis, the mirror plane $\sigma_{v}$ is assumed in order as the $zx$- and $yz$-plane, and the mirror plane $\sigma_{h}$ is assumed as the $x$-$y$ plane. We use the MSGCorep codes Liu et al. (2023b) to filter the groups without time-reversal and inversion but with the combined time-reversal and inversion symmetry and find their matrices.
Groups This work QMD
$\begin{matrix}&\bar{1}^{\prime},~{}2^{\prime}/\text{m}\end{matrix}$ $\begin{matrix}xxx&xyy\\ yxx&yyy\end{matrix}$ $\begin{matrix}yxx&xyy\end{matrix}$
$\begin{matrix}&2/\text{m}^{\prime},~{}\text{m}^{\prime}\text{m}\text{m}\end{matrix}$ $\begin{matrix}yxx&yyy\end{matrix}$ $\begin{matrix}yxx\end{matrix}$
$\begin{matrix}&\bar{3}^{\prime}\text{m},\bar{3}^{\prime}\text{m}^{\prime},~{}6^{\prime}/\text{m}\text{m}\text{m}^{\prime}\end{matrix}$ $\begin{matrix}yxx=-yyy\end{matrix}$$\times$
$\begin{matrix}&\bar{3}^{\prime},~{}6^{\prime}/\text{m}\end{matrix}$ $\begin{matrix}xyy=-xxx\\ yxx=-yyy\end{matrix}$$\times$
[…]
Nonzero quadratic conductivity elements at quantum metric order for magnetic point groups.– With the help of the above formulas, we can determine the non-zero quadratic nonlinear conductivity elements for the magnetic groups to facilitate experimental verification (details in Sec. SIIIB of Sup ). At the order of quantum metric, the system has to break inversion and time-reversal symmetries but preserve the combined time-reversal and inversion symmetry, narrowing the candidates to only 21 out of a total of 122 magnetic point groups. Take the nonzero elements on the $x$-$y$ plane for the $\bar{3}^{\prime}m^{\prime}$ groups for example. One of the symmetry operators is $C_{3}\mathcal{P}\mathcal{T}$, which can be expressed as a matrix $\mathcal{R}$ that rotates the Cartesian coordinates and a matrix $\mathcal{Q}$ that acts on the components of the quadratic electric fields $\{E_{x}^{2},2E_{x}E_{y},E_{y}^{2}\}$. Applying the $C_{3}\mathcal{P}\mathcal{T}$ operation to $J_{a}=\chi_{abb}E_{b}E_{b}$ is equivalent to $\chi=\mathcal{R}\chi\mathcal{Q}^{-1}$, from which the relations $\chi_{yxx}=-\chi_{yyy}$ and $\chi_{xyy}=-\chi_{xxx}$ can be found, as well as $\chi_{abb}^{\mathrm{QMD}}=0$ because the anti-symmetric part $\chi_{aaa}^{\mathrm{QMD}}=0$ (details in Sec. SIIIB of Sup ). The procedure is applied to all symmetry operations in the $\bar{3}^{\prime}m^{\prime}$ group, and the common relations give rise to the nonzero quadratic conductivity elements $\chi_{abb}$ on the $x$-$y$ plane ($\{a,b\}\in\{x,y\}$) as well as their relations, as shown in Tab. 2 for only 9 out of the 122 magnetic point groups. Only 20 magnetic point groups have nonzero quadratic conductivity elements in 3D ($\chi_{abc}$, $a,b,c\in\{x,y,z\}$) (see Tabs. S6-S12 of Sup ).
MnBi2Te4 films with and without $C_{3}$ symmetry.– To test our formulas, we consider MnBi2Te4 thin films with and without the $C_{3}$ symmetry (Sec. SIV of Sup ), because the $C_{3}$ symmetry forbids conductivity on the $x$-$y$ plane with anti-symmetric labels, i.e., $\chi_{abc}=-\chi_{cba}$, ($(\{a,b,c\}\in\{x,y\})$ Du et al. (2021b). The minimal Hamiltonian reads
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{0}+\hat{\mathcal{H}}_{S},$ (4)
where the $C_{3}$ symmetric part Sun et al. (2020) ^$\hat{\mathcal{H}}_{0}=h_{0}+h_{k}\hat{\tau}_{z}\hat{\sigma}_{z}-\gamma k_{x}\hat{\tau}_{0}\hat{\sigma}_{y}+\gamma k_{y}\hat{\tau}_{0}\hat{\sigma}_{x}+m_{k}\hat{\tau}_{x}\hat{\sigma}_{z}$, the Pauli matrices $\hat{\tau}$ and $\hat{\sigma}$ describe orbital and spin degrees of freedom, respectively, $h_{0}=E_{0}-Dk^{2}$, $k_{x}$ and $k_{y}$ are wave vectors, $k^{2}\equiv k_{x}^{2}+k_{y}^{2}$, $h_{k}=\Delta/2-Bk^{2}$, $m_{k}=m_{2}+\beta_{1}\left(3k_{x}^{2}k_{y}-k_{y}^{3}\right)$, $E_{0}$, $D$, $\Delta$, $B$, $m_{2}$, $\beta_{1}$, and $\gamma$ are model parameters. The $C_{3}$-symmetry-breaking part reads Gao et al. (2023) ^$\hat{\mathcal{H}}_{S}=(s_{1}+s_{3})k_{y}^{2}\hat{\tau}_{0}\hat{\sigma}_{0}/2+(s_{1}-s_{3})k_{y}^{2}\hat{\tau}_{z}\hat{\sigma}_{z}/2-s_{5}k_{y}\hat{\tau}_{0}\hat{\sigma}_{y}+s_{7}k_{y}\hat{\tau}_{x}\hat{\sigma}_{z}+s_{8}k_{x}\hat{\tau}_{y}\hat{\sigma}_{z}$, where $s_{1}$, $s_{3}$, $s_{5}$, $s_{7}$, and $s_{8}$ are model parameters. The model is also classified into the $\bar{3}^{\prime}/m^{\prime}$ and $2/m^{\prime}$ magnetic point groups, respectively, as shown by the Fermi surfaces and energy bands in Fig. 2 (a1)-(b2). Figure 2 (c1)-(j2) shows the total and splits of contributions to the quadratic nonlinear conductivity at the order of quantum metric as functions of the Fermi energy, by applying the formulas in Tab. 1 to the model in Eq. (4). Several features verify the formulas. (i) The $C_{3}$ symmetry cancels $\chi^{\mathrm{QMD}}$ Wang et al. (2021); Liu et al. (2021), which can be recovered by the strain that breaks the $C_{3}$ symmetry Gao et al. (2023). Our BSJ and BSK also vanish (emerge) under (without) the $C_{3}$ symmetry, consistent with the analysis in Tab. 2. (ii) The contributions of QMD, BSJ, BSK are comparable, but 2SJ, 2SK, and SSK could overwhelm QMD by orders, when the parameter
$\lambda\equiv n_{i}V_{1}^{3}/(n_{i}V_{0}^{2})^{2}\propto 1/n_{i}$ (5)
is large ($n_{i}$ is small, i.e., the clean limit), where $n_{i}$ is the impurity density, $V_{0}^{2}$ ($V_{1}^{3}$) measures the correlations of two (three) scattering events. $\chi^{\mathrm{QMD}}$, $\chi^{\mathrm{BSJ}}$, $\chi^{\mathrm{BSK}}_{2}$, and $\chi^{\mathrm{2SJ}}_{1,2,3}$ do not depend on $\lambda$, $\chi^{\mathrm{BSK}}_{1}$ is linear in $\lambda$, $\chi^{\mathrm{SSK}}_{1,2,3,4}$ has terms $\propto\lambda^{0}$ and $\lambda^{1}$, $\chi^{\mathrm{2SK}}_{1,2}$ has terms $\propto\lambda^{0}$, $\lambda^{1}$, and $\lambda^{2}$ (Tab. S5 of Sup ). The Drude contribution Sodemann and Fu (2015); Wang et al. (2021) depends on $1/(n_{i}V_{0}^{2})^{2}$. Also, for $\chi_{yxy}$ of a specific model, the disorder-induced but disorder-strength-independent contributions could overwhelms the quantum metric part Atencia et al. (2023).
Figure 3: Scaling laws between the quadratic Hall conductivity $\chi_{yxx}$ and linear longitudinal conductivity $\sigma_{xx}$ for the experimental data (circles) of (a) 4SL Wang et al. (2023) and (b) BP-6SL Gao et al. (2023) MnBi2Te4 thin films (SL for septuple layer, BP for black phosphorous), fitted using Eqs. (6) (solid) and (7) (dashed, details in Fig. S3 of Sup ). The fitting parameters in (a) $\mathcal{C}_{2}=5.05\times 10^{4}~{}\mathrm{mA\cdot nm}$, $\mathcal{C}_{1}=-1.61\times 10^{4}~{}\mathrm{nm\cdot V^{-1}}$, $\mathcal{C}_{0}=1.03\times 10^{3}~{}\mathrm{mA\cdot nm\cdot V^{-1}}$, and in (b) $\mathcal{C}_{2}=6.98$, $\mathcal{C}_{1}=-3.38$, $\mathcal{C}_{0}=1.05\times 10^{2}$, with the same units as those in (a). The data in (b) is multiplied by 8 nm to be treated as a 2D case.
Scaling law at the order of quantum metric.– The scaling law is powerful in distinguishing mechanisms in the Hall effects Nagaosa et al. (2010); Tian et al. (2009); Hou et al. (2015); Du et al. (2019, 2021a); Tiwari et al. (2021); Duan et al. (2022). The formulas in Tab. 1 allow us to extend the scaling law of the nonlinear Hall effect to the order of quantum metric (details in Sec. SV of Sup ), characterized by the relation between the nonlinear Hall conductivity $\chi_{yxx}$ and longitudinal linear conductivity $\sigma_{xx}$ as
$\chi_{yxx}=\mathcal{C}_{2}\sigma_{xx}^{2}+\mathcal{C}_{1}\sigma_{xx}+\mathcal{C}_{0},$ (6)
where in experiments $\chi_{yxx}=[V^{2\omega}_{y}\sigma_{xx}/(V^{\omega}_{x})^{2}](L^{2}/W)$, with $L$ the length, $W$ the width in 2D or cross-sectional area in 3D, as well as the transverse double-frequency voltage $V^{2\omega}_{y}$ and longitudinal single-frequency voltage $V^{\omega}_{x}$. The 3 fitting parameters $\mathcal{C}_{a}$ have physical meanings, as categorized in Tab. 3.
Table 3: Classification of the fitting parameters in the scaling law Eq. (6) by the mechanisms in Fig. 1 and Tab. 1 and their dependence on the impurity density $n_{i}$. The superscripts $(3A)$ and $(4A)$ indicate how many times the anti-symmetric scattering rates $\varpi_{ll^{\prime}}^{(3A)}$ $\sim$ $n_{i}V_{1}^{3}$ and $\varpi_{ll^{\prime}}^{(4A)}$ $\sim$ $(n_{i}V_{0}^{2})^{2}$ from $\varpi_{ll^{\prime}}^{A}$ appear in the formulas of $\chi^{\mathrm{2SK}}_{1,2,3,4}$ and $\chi^{\mathrm{SSK}}_{1,2}$.
Param. Mechanisms
$\mathcal{C}_{2}\sim n_{i}^{-2}$ $\chi_{1,2}^{2\mathrm{SK}(3A,3A)},\chi^{\mathrm{Drude}}$
$\mathcal{C}_{1}\sim n_{i}^{-1}$ $\chi_{1}^{\mathrm{BSK}},\chi_{1,2}^{2\mathrm{SK}(3A,4A)},\chi_{1,2,3,4}^{\mathrm{SSK}(3A)}$
$\mathcal{C}_{0}\sim n_{i}^{0}$ $\chi_{2}^{\mathrm{BSK}},\chi^{\mathrm{BSJ}},\chi^{2\mathrm{SJ}}_{1,2,3},\chi_{1,2}^{2\mathrm{SK}(4A,4A)},\chi_{1,2,3,4}^{\mathrm{SSK}(4A)},\chi^{\mathrm{QMD}}$
The fitting in Fig. 3 with Eq. (6) shows that all terms are competing from 2 to 10 K in 4SL-layered MnBi2Te4 Wang et al. (2023), e.g., at $\sigma_{xx}=170\ \mu$S, $\mathcal{C}_{2}\sigma_{xx}^{2}=1.46\times 10^{3}$, $\mathcal{C}_{1}\sigma_{xx}^{1}=-2.73\times 10^{3}$, and $\mathcal{C}_{0}=1.03\times 10^{3}$; and from 2 to 15 K in BP-6SL-MnBi2Te4 thin films Gao et al. (2023), the $\mathcal{C}_{0}$ term is dominant, e.g., at $\sigma_{xx}=540\ \mu$S, $\mathcal{C}_{2}\sigma_{xx}^{2}=2.04$, $\mathcal{C}_{1}\sigma_{xx}^{1}=-1.84$, and $\mathcal{C}_{0}=105$. Nevertheless, all mechanisms could contribute in $\mathcal{C}_{0}$. The scaling law Eq. (6) considers only the static scattering. If the dynamic scattering (e.g., due to phonons) is also included, the scaling law becomes
$\chi_{yxx}=\mathcal{C}_{4}^{\prime}\sigma_{xx}^{4}+\mathcal{C}_{3}^{\prime}\sigma_{xx}^{3}+\mathcal{C}_{2}^{\prime}\sigma_{xx}^{2}+\mathcal{C}_{1}^{\prime}\sigma_{xx}+\mathcal{C}_{0}^{\prime},$ (7)
and the fitting results show more detailed and entangled origins of mechanism (see Fig. S3 of Sup ). A similar scaling law is found Huang et al. (2023), but no explicit conductivity formulas are given as those in Tab. 1 and the parameters are not classified according to the physical mechanisms in Fig. 1 and Tab. 3.
Acknowledgements.
Z. H. Gong and Z. Z. Du contributed equally. This work was supported by the National Key R&\&&D Program of China (2022YFA1403700), Innovation Program for Quantum Science and Technology (2021ZD0302400), the National Natural Science Foundation of China (11925402, 12374041, 12350402), Guangdong Basic and Applied Basic Research Foundation (2023B0303000011), Guangdong province (2020KCXTD001), the Science, Technology and Innovation Commission of Shenzhen Municipality (ZDSYS20190902092905285), and Center for Computational Science and Engineering of SUSTech.